40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 6 01:09:11 2017, johannes, Update, Computer Scripts / Programs, New set of loss measurements XARM_20170905.pdfYARM_20170905.pdf
    Reply  Mon Sep 11 12:56:40 2017, johannes, Update, Computer Scripts / Programs, lossmap attempts 
       Reply  Mon Sep 11 16:46:09 2017, Steve, Update, PEM, earthquakes M5.4eq.png
          Reply  Tue Sep 19 14:00:41 2017, Steve, Update, PEM, earthquakes 7.1MX5.3J.png
             Reply  Fri Oct 27 09:34:20 2017, Steve, Update, PEM, earthquakes recentEQs.png
                Reply  Fri Nov 10 13:46:26 2017, Steve, Update, PEM, M3.1 local earthquake 3.1M_local_EQ.png
                   Reply  Mon Apr 9 08:39:39 2018, Steve, Update, PEM, M5.3 eq Souther CA M5.3_Santa_Cruz_Is.CA.pngafter_M5.3.pngM5.3vac.png
                Reply  Thu Jan 25 08:27:25 2018, Steve, Update, PEM, earthquakes M4_local_eq.pngM4.pngM5.8Ferndale_CA.png
       Reply  Mon Sep 25 17:56:33 2017, johannes, Update, Computer Scripts / Programs, transmitted power during lossmap 
    Reply  Thu Oct 11 13:40:51 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
       Reply  Fri Oct 12 12:29:34 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
          Reply  Fri Oct 12 20:20:29 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
             Reply  Sat Oct 13 20:11:10 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
                Reply  Mon Oct 15 10:32:13 2018, yuki, Update, Computer Scripts / Programs, loss measurements XARM_20181015_1500.pdfYARM_20181015_1500.pdf
                   Reply  Tue Oct 16 00:44:29 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
                   Reply  Fri Nov 2 19:25:16 2018, gautam, Update, Computer Scripts / Programs, loss measurements 
    Reply  Mon Oct 15 12:52:54 2018, yuki, Update, Computer Scripts / Programs, additional comments 
       Reply  Fri Nov 2 16:42:31 2018, aaron, Update, Computer Scripts / Programs, arm loss measuremenents 
          Reply  Mon Nov 5 13:52:18 2018, aaron, Update, Computer Scripts / Programs, arm loss measuremenents 
             Reply  Tue Nov 6 10:19:26 2018, aaron, Update, Computer Scripts / Programs, arm loss measuremenents 
                Reply  Tue Nov 6 15:23:48 2018, gautam, Update, IOO, IMC problematic MCwonky.png
                   Reply  Tue Nov 6 19:02:35 2018, aaron, Update, IOO, IMC problematic 181106_MCTRANS.jpg
                      Reply  Sat Nov 10 17:40:00 2018, aaron, Update, IOO, IMC problematic 
                         Reply  Mon Nov 12 13:53:20 2018, rana, Update, IOO, loss measurement: oscope vs CDS DAQ 
                         Reply  Thu Nov 15 10:21:07 2018, aaron, Update, IOO, IMC problematic 181115_chansDown.pngPD_noise.png
                            Reply  Fri Nov 16 10:53:07 2018, aaron, Update, IOO, IMC problematic Screenshot_from_2018-11-16_19-22-34.png
                               Reply  Sat Nov 17 18:59:01 2018, aaron, Update, IOO, IMC problematic 
                Reply  Wed Nov 7 05:16:16 2018, yuki, Update, Computer Scripts / Programs, arm loss measuremenents 
Message ID: 14243     Entry time: Thu Oct 11 13:40:51 2018     In reply to: 13299     Reply to this: 14245
Author: yuki 
Type: Update 
Category: Computer Scripts / Programs 
Subject: loss measurements 
Quote:

This is the procedure I follow when I take these measurements for the XARM (symmetric under XARM <-> YARM):

  1. Dither-align the interferometer with both arms locked. Freeze outputs when done.
  2. Misalign ETMY + ITMY.
  3. ITMY needs to be misaligned further. Moving the slider by at least +0.2 is plentiful to not have the other beam interfere with the measurement.
  4. Start the script, which does the following:
    1. Resume dithering of the XARM
    2. Check XARM dither error signal rms with CDS. If they're calm enough, proceed.
    3. Freeze dithering
    4. Start a new set of averages on the scope, wait T_WAIT (5 seconds)
    5. Read data (= ASDC power and MC2 trans) from scope and save
    6. Misalign ETMX and wait 5s
    7. Read data from scope and save
    8. Repeat desired amount of times
  5. Close the PSL shutter and measure the PD dark levels

Information for the armloss measurement:

  • Script which gets the data:  /users/johannes/40m/armloss/scripts/armloss_scope/armloss_dcrefl_asdcpd_scope.py
  • Script which calculates the loss: /users/johannes/40m/armloss/scripts/misc/armloss_AS_calc.py
  • Before doing the procedure Johannes wrote you have to prepare as follows:
    • put a PD in anti-symmetric beam path to get ASDC signal.
    • put a PD in MC2 box to get tranmitted light of IMC. It is used to normalize the beam power.
    • connect those 2 PDs to oscilloscope and insert an internet cable to it.
  • Usage: python2 armloss_dcrefl_asdcpd_scope.py [IP address of Scope] [ScopeCH for AS] [ScopeCH for MC] [Num of iteration] [ArmMode]

Note: The scripts uses httplib2 module. You have to install it if you don't have.

The locked arms are needed to calculate armloss but the alignment of PMC is deadly bad now. So at first I will make it aligned. (Gautam aligned it and PMC is locked now.) 

gautam: The PMC alignment was fine, the problem was that the c1psl slow machine had become unresponsive, which prevented the PMC length servo from functioning correctly. I rebooted the machine and undid the alignment changes Yuki had made on the PSL table.

ELOG V3.1.3-